# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v1

@@ 2.1 The shape of order (anti)isomorphisms
We start with a simple result showing that it is enough to study the order isomorphisms rather than the more arduous order anti-isomorphisms or anti-involutions.
@@
If $\mathcal{G}=\operatorname*{Rg}(B)$ as in (10), then, for every $f\in\operatorname*{Rg}(B)$, there exists $(w_{y})_{y\in\mathscr{X}}$ such that $f(\cdot)=\sup_{y\in\mathscr{X}}w_{y}+b(\cdot,y)$. Hence we have that $J(f)=\sup_{y\in\mathscr{X}}J(w_{y}+b(\cdot,y))(\cdot)$ as $J$ commutes with the $\sup$, as shown in (9). Thus to characterize $J$, we just have to study its action on the functions $\{w+b(\cdot,y)\}_{y\in\mathscr{X},w\in\mathbb{R}_{\bot}}$. More generally this argument holds for sup-generators. However, unless there are other assumptions, such as continuity or compacity, see 26, the $\{w+b(\cdot,y)\}_{y\in\mathscr{X},w\in\mathbb{R}_{\bot}}$ are not necessarily sup-irreducible or relatively-inf-irreducible elements, which have the nice property of being stable under order isomorphisms.
@@
Every order isomorphism $J:\mathscr{F}\rightarrow\mathcal{G}$ over sets ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ sends sup-irreducible elements (resp. $\mathscr{F}$-relatively-inf-irreducible) of $\mathscr{F}$ onto sup-irreducible elements (resp. $\mathcal{G}$-relatively-inf-irreducible) of $\mathcal{G}$.
Every anti-involution $T$ over $\mathcal{G}$ sends $\mathcal{G}$-relatively-inf-irreducible elements of $\mathcal{G}$ onto sup-irreducible elements of $\mathcal{G}$.
@@
Let $f_{0}\in\mathscr{F}$ be sup-irreducible in $\mathscr{F}$. Take $g_{1},h_{1}\in\mathcal{G}$ such that $J(f_{0})=\sup(g_{1},h_{1})$. Since $J$ is onto, we can find $g_{0},h_{0}\in\mathscr{F}$ such that $g_{1}=J(g_{0})$, $h_{1}=J(h_{0})$. Hence, using that $J$ commutes with the $\sup$, $J(f_{0})=J(\sup(g_{0},h_{0}))$. Since $J$ is bijective and $f_{0}$ is sup-irreducible, we have that $f_{0}=g_{0}$ or $f_{0}=h_{0}$, applying $J$, $J(f_{0})$ is thus sup-irreducible in $\mathscr{F}$. For $\mathscr{F}$-relatively-inf-irreducible functions we have similarly that if $J(f_{0})=\inf\nolimits^{\mathcal{G}}(g_{1},h_{1})$, then $f_{0}=\sup\{J^{-1}h\,|\,h\in\mathcal{G},\,h\leq g_{1},\,h\leq h_{1}\}$, so $f_{0}=\inf\nolimits^{\mathscr{F}}(J^{-1}g_{1},J^{-1}h_{1})$. The statement concerning anti-involutions reduces to the case of order isomorphisms, since an anti-involution becomes an order isomorphism, when the origin or destination set is equipped with the opposite order. ∎
@@
In [AAFM12, Definitions 4.14,4.15], the authors say that a function $f\in\mathcal{G}$ is $P$-extremal if, for all $g,h\in\mathcal{G}$, $(f\leq g)\&(f\leq h)\Rightarrow g\leq h\text{ or }h\geq g$, and conversely $Q$-extremal if, for all $g,h\in\mathcal{G}$, $(f\geq g)\&(f\geq h)\Rightarrow g\leq h\text{ or }h\geq g$. In other words the upper/lower bounds on an extremal $f$ should be “comparable”. With the same proof as that of 13, one can show that $P$ and $Q$-extremal functions are preserved by order isomorphisms. Moreover, if $f$ is $P$-extremal, then $f=\inf\nolimits^{\mathcal{G}}(g,h)\Rightarrow f\leq g,h\stackrel{{\scriptstyle w.l.o.g}}{{\Rightarrow}}g\leq h\Rightarrow f=g$, so $f$ is relatively-inf-irreducible. For some sets, such as lower semicontinuous convex functions on a vector space, $P$-extremal functions and relatively-inf-irreducible ones coincide, since they are the “Dirac masses” $\delta^{\top}_{x}+\lambda$. However, for $1$-Lipschitz functions, there are simply no $P$-extremal functions, [^1] This is shown for instance by taking two points $x_{0}\neq x_{1}$. Then any $1$-Lipschitz function $f$ satisfies $f\leq\inf(d(\cdot,x_{0})+\lambda_{0},d(\cdot,x_{1})+\lambda_{0})$ with $\lambda_{0}=\max(f(x_{0}),f(x_{1}))$, but the two upper bounds are not comparable. whereas the distance functions $d(\cdot,x_{0})+\lambda$ are all relatively-inf-irreducible. This is why we favor the concept of irreducibility over comparability, even though the latter proved fruitful in [AAFM12, Section 4].
Sometimes, we may want to study sets that are not sup-stable, such as that of continuous functions or of Lipschitz functions, both of which are pointwise dense in the sup-stable set of lower semicontinuous functions. We can however extend max-isomorphisms over a set to the smallest sup-stable set containing it. The proof of the following is inspired by the theory of Galois connections, see [${}^{+}$03, Chapter 0, Section 3].
@@
Let $J:\mathscr{F}\rightarrow\mathcal{G}$ be an order isomorphism between some non-empty sets ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$, such that $\mathscr{F}$ (resp. $\mathcal{G}$) is pointwise dense in the sup-closure ¯supsuperscriptnormal-¯$\overline{\mathscr{F}}^{\sup}$ of $\mathscr{F}$ (resp. $\overline{\mathcal{G}}^{\sup}$). Then $J$ can be extended to an order isomorphism between ¯supsuperscriptnormal-¯$\overline{\mathscr{F}}^{\sup}$ and $\overline{\mathcal{G}}^{\sup}$. The density holds in particular if the sets are finitely sup-stable, i.e. if $f,g\in\mathscr{F}$ (resp. $\mathcal{G}$), then $\sup(f,g)\in\mathscr{F}$ (resp. $\mathcal{G}$).
@@
where $J:\mathscr{F}\rightarrow\mathcal{G}$ is an order isomorphism. The irreducible elements of $\mathscr{F}^{\prime}$ are simply those of $\mathscr{F}$ translated by $f_{0}$. The results above also hold when “order” is replaced with (max,+).
@@
$\bar{J}f=\sup\{J(g)\,|\,g\in\mathscr{F},\,g\leq f\},$ (13)
$\bar{J}^{1}g=\sup\{J^{-1}(g)\,|\,g\in\mathcal{G},\,g\leq f\}.$ (14)
Clearly their restrictions to $\mathscr{F}$ and $\mathcal{G}$ are equal to $J$ and $J^{-1}$ respectively and are order preserving. We want to show that $f=\bar{J}^{1}\bar{J}f$. Fix $f\in\overline{\mathscr{F}}^{\sup}$, and set $\tilde{f}=\bar{J}^{1}\bar{J}f$. We then have that
@@
Fix $h\in\mathscr{F}$ such that $J(h)\leq\sup_{h^{\prime}\leq f}J(h^{\prime})$, hence $J(h)\in\{g^{\prime}\in\mathcal{G}\,|\,J^{-1}(g^{\prime})\leq f\}$, so $h\leq f$. Taking the supremum over $h$ shows that $\tilde{f}\leq f$.
To prove the converse, we are going to show that, for any $f\in\overline{\mathscr{F}}^{\sup}$ and $g\in\overline{\mathcal{G}}^{\sup}$, we have that $g\geq\bar{J}f\implies\bar{J}^{1}g\geq f$. Take any $f,g$ satisfying $g\geq\bar{J}f$. Fix $f^{\prime}\in\mathscr{F}$ such that $f^{\prime}\leq f$, so $Jf^{\prime}\leq\bar{J}f\leq g$, whence $f^{\prime}\leq\bar{J}^{1}g$. Taking the supremum over $f^{\prime}$, since $\mathscr{F}$ was assumed dense in ¯supsuperscript¯$\overline{\mathscr{F}}^{\sup}$ for the pointwise topology, we obtain $\bar{J}^{1}g\geq f$. Since $\bar{J}f\geq\bar{J}f$, the implication we have just shown entails that $\bar{J}^{1}\bar{J}f\geq f$. We thus obtained that $f=\bar{J}^{1}\bar{J}f$, so $\bar{J}$ is an order isomorphism.
We now show that $\mathscr{F}$ is dense in $\overline{\mathscr{F}}$ for the pointwise topology if $\mathscr{F}$ is finitely sup-stable. Indeed, let $f\in\overline{\mathscr{F}}^{\sup}$ and fix $(f_{\alpha})_{\alpha\in\mathcal{A}}$ such that $f=\sup_{\alpha}f_{\alpha}$. Denote by $\mathcal{B}$ the set of finite subsets of $\mathcal{A}$, and for $B\in\mathcal{B}$, set $f_{B}=\sup_{\alpha\in B}f_{\alpha}$. Then $f=\sup_{B\in\mathcal{B}}f_{B}$, e.g. by taking $B=\{\alpha\}$, however $\sup_{B\in\mathcal{B}}f_{B}=\lim_{B\in\mathcal{B}}\uparrow f_{B}$ as the pointwise limit of the net. This proves the density result.
For (12), just set $J(f-f_{0}):=J^{\prime}(f)-g_{0}$ with inverse $J^{-1}(g-g_{0})=(J^{\prime})^{-1}(f)-f_{0}$. Take a sup-irreducible $f\in\mathscr{F}$, then, for all $g,h\in\mathscr{F}$, $f+f_{0}=\sup(g+f_{0},h+f_{0})\implies f=g\text{ or }f=h$, so $(f+f_{0})$ is sup-irreducible in $\mathscr{F}^{\prime}$. The other implications for irreducible functions are shown similarly. ∎
It was rightly mentioned in [AAM09, p670] that most results could be obtained either starting from order preserving or order reversing operators, without much difference. However we will focus on order preserving ones, as done in the convex case by [IRS15, p75]. Indeed order preserving operators have the advantage that the identity operator is always order preserving, whereas we will meet spaces with no anti-involution. Most of the results stem from characterizing either the sup or inf-irreducible elements. The usefulness of focusing on such elements was exemplified in [GK07, p5] for finite sets $\mathscr{X}$.

### v2

@@ 2.1 The shape of order (anti)isomorphisms
We start with a simple result showing that, if we know one reference anti-isomorphism, as given for instance by 10, it is enough to study the order isomorphisms rather than the more arduous order anti-isomorphisms or anti-involutions.
@@
If $\mathcal{G}=\operatorname*{Rg}(B)$ as in (10), then, for every $f\in\operatorname*{Rg}(B)$, there exists $(w_{y})_{y\in\mathscr{Y}}$ such that $f(\cdot)=\sup_{y\in\mathscr{Y}}w_{y}+b(\cdot,y)$. Hence we have that $J(f)=\sup_{y\in\mathscr{Y}}J(w_{y}+b(\cdot,y))(\cdot)$ as $J$ commutes with the $\sup$, as shown in (9). Thus to characterize $J$, we just have to study its action on the functions $\{w+b(\cdot,y)\}_{y\in\mathscr{T},w\in\mathbb{R}_{\bot}}$. More generally this argument holds for sup-generators. However, unless there are other assumptions, such as continuity or compactness, see 26, the $\{w+b(\cdot,y)\}_{y\in\mathscr{T},w\in\mathbb{R}_{\bot}}$ are not necessarily sup-irreducible or relatively-inf-irreducible elements, which have the nice property of being stable under order isomorphisms.
Let us now state how irreducible elements are invariants of order isomorphisms.
@@
Every order isomorphism $J:\mathscr{F}\rightarrow\mathcal{G}$ over sup-stable sets ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ sends sup-irreducible elements (resp. $\mathscr{F}$-relatively-inf-irreducible) of $\mathscr{F}$ onto sup-irreducible elements (resp. $\mathcal{G}$-relatively-inf-irreducible) of $\mathcal{G}$.
Every anti-isomorphism $T:\mathscr{F}\to\mathcal{G}$ sends $\mathscr{F}$-relatively-inf-irreducible elements of $\mathscr{F}$ onto sup-irreducible elements of $\mathcal{G}$.
@@
We just apply 7. Let $f_{0}\in\mathscr{F}$ be sup-irreducible in $\mathscr{F}$. Take $g_{1},h_{1}\in\mathcal{G}$ such that $Jf_{0}=\sup(g_{1},h_{1})$. Since $J$ is onto, we can find $g_{0},h_{0}\in\mathscr{F}$ such that $g_{1}=Jg_{0}$, $h_{1}=Jh_{0}$. Hence, using that $J$ commutes with the $\sup$, $Jf_{0}=J(\sup(g_{0},h_{0}))$. Since $J$ is bijective and $f_{0}$ is sup-irreducible, we have that $f_{0}=g_{0}$ or $f_{0}=h_{0}$, applying $J$, $Jf_{0}$ is thus sup-irreducible in $\mathcal{G}$. For $\mathscr{F}$-relatively-inf-irreducible functions $f_{0}$ we have similarly that if $Jf_{0}=\inf\nolimits^{\mathcal{G}}(g_{1},h_{1})$, then $f_{0}=\sup\{J^{-1}h\,|\,h\in\mathcal{G},\,h\leq g_{1},\,h\leq h_{1}\}$, so $f_{0}=\inf\nolimits^{\mathscr{F}}(J^{-1}g_{1},J^{-1}h_{1})$.
Concerning $T:\mathscr{F}\to\mathcal{G}$, if $Tf=\sup(g_{1},h_{1})$, then, by 7, $f=\inf\nolimits^{\mathscr{F}}(T^{-1}g_{1},T^{-1}h_{1})$, and we conclude as before. ∎
@@
In [AAFM12, Definitions 4.14,4.15], the authors say that a function $f\in\mathcal{G}$ is $P$-extremal if, for all $g,h\in\mathcal{G}$, $(f\leq g)\&(f\leq h)\Rightarrow g\leq h\text{ or }h\geq g$, and conversely $Q$-extremal if, for all $g,h\in\mathcal{G}$, $(f\geq g)\&(f\geq h)\Rightarrow g\leq h\text{ or }h\geq g$. In other words the upper/lower bounds on an extremal $f$ should be “comparable”. With the same proof as that of 13, one can show that $P$ and $Q$-extremal functions are preserved by order isomorphisms. Moreover, if $f$ is $P$-extremal, then $f=\inf\nolimits^{\mathcal{G}}(g,h)\Rightarrow f\leq g,h\stackrel{{\scriptstyle w.l.o.g}}{{\Rightarrow}}g\leq h\Rightarrow f=g$, so $f$ is relatively-inf-irreducible. For some sets, such as that of lower semicontinuous convex functions on a vector space, $P$-extremal functions and relatively-inf-irreducible ones coincide, since they are the “Dirac masses” $\delta^{\top}_{x}+\lambda$. However, for $1$-Lipschitz functions, there are simply no $P$-extremal functions, [^3] This is shown for instance by taking two points $x_{0}\neq x_{1}$. Then any $1$-Lipschitz function $f$ satisfies $f\leq\inf(d(\cdot,x_{0})+\lambda_{0},d(\cdot,x_{1})+\lambda_{0})$ with $\lambda_{0}=\max(f(x_{0}),f(x_{1}))$, but the two upper bounds are not comparable. whereas the distance functions $d(\cdot,x_{0})+\lambda$ are all relatively-inf-irreducible. This is why we favor the concept of irreducibility over comparability, even though the latter proved fruitful in [AAFM12, Section 4].
Sometimes, we may want to study sets that are not sup-stable, such as that of continuous functions or of Lipschitz functions, both of which are pointwise dense in the sup-stable set of lower semicontinuous functions. Under some regularity assumptions, we can however extend order isomorphisms over a set to the smallest sup-stable set containing it. The proof of the following is inspired by the theory of Galois connections, see [GHK+03, Chapter 0, Section 3].
@@
Let $J:\mathscr{F}\rightarrow\mathcal{G}$ be an order isomorphism between some non-empty sets ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ stable by addition of scalars and finitely sup-stable. Assume also that
$\forall g\in\mathcal{G},\,\forall g^{\prime}=\lim_{\alpha}\uparrow g_{\alpha}\in\overline{\mathcal{G}}^{\sup}\text{ with $h_{\alpha}\in\mathcal{G}$, }(g\leq g^{\prime})\implies(\forall\epsilon>0,\,\exists\alpha_{\epsilon}\text{ s.t.\ }g\leq g_{\alpha_{\epsilon}}+\epsilon),$ (12)
where $\lim_{\alpha}\uparrow g_{\alpha}$ denotes a pointwise non-decreasing net, and suppose that
$\forall g^{\prime}\in\overline{\mathcal{G}}^{\sup},\,\forall y\in\mathscr{Y},\,\exists v_{y}:(0,+\infty)\to\mathbb{R}\text{ s.t.\ }\lim_{\epsilon^{\prime}\to 0}v_{y}(\epsilon^{\prime})=0\text{ and }\\ \forall g\in\mathcal{G},(g\leq g^{\prime})\implies\left(\forall\epsilon>0,\,J^{-1}(g+\epsilon)(y)\leq J^{-1}g(y)+v_{y}(\epsilon)\right).$ (13)
Then $J$ can be extended to an order isomorphism between ¯supsuperscript¯$\overline{\mathscr{F}}^{\sup}$ and $\overline{\mathcal{G}}^{\sup}$.
@@
where $J:\mathscr{F}\rightarrow\mathcal{G}$ is an order isomorphism. The irreducible elements of $\mathscr{F}^{\prime}$ are simply those of $\mathscr{F}$ translated by $f_{0}$. The results above also hold when “order” is replaced with (max,+), in particular (13) is then automatically satisfied with $v_{y}(\epsilon)=\epsilon$.
For instance (12) is satisfied if we have that $\mathcal{G}$ is a subset of the continuous functions on a compact set $\mathscr{X}$. Indeed we then have that for any fixed $\epsilon>0$
$\forall x\in\mathscr{X},\exists\alpha_{x}\text{ s.t.\ }\epsilon+g_{\alpha_{x}}(x)\geq\min(g^{\prime}(x),g(x)),\text{ and }\\ \exists V_{x,\alpha_{x}}^{g}\text{ open neighborhood of $x$},\,\forall y\in V_{x,\alpha_{x}}^{g},\,\max(|g(x)-g(y)|,|g_{\alpha_{x}}(x)-g_{\alpha_{x}}(y)|)\leq\epsilon$
In other words: $\forall x\in\mathscr{X},\exists\alpha_{x},\,g\leq g_{\alpha_{x}}+3\epsilon$. Now since $\mathscr{X}\subset\bigcup_{x\in\mathscr{X}}V_{x,\alpha_{x}}^{g}$ and $\mathscr{X}$ is assumed compact, we can find a finite subcover indexed by $\{x_{1},\dots,x_{N}\}$ and $g\leq\sup_{i}g_{\alpha_{x_{i}}}+3\epsilon$ over $\mathscr{X}$. Since $g_{\alpha}$ was assumed to be increasing there is a $\bar{\alpha}$ such that $\sup_{i}g_{\alpha_{x_{i}}}\leq g_{\bar{\alpha}}$. Alternatively it can be seen as a consequence of Dini’s theorem applied to $g^{\prime}_{\alpha}=\sup(g,g_{\alpha})$. Eq.(12) still holds if $\mathscr{X}$ is no longer compact but the functions “tend to 00 at infinity”, i.e. for all $h\in\mathcal{G}$, $\epsilon>0$, there exists a compact $K_{\epsilon}\subset\mathscr{X}$ s.t. for $x\notin K_{\epsilon}$, $|h(x)|\leq\epsilon$.
@@
$\bar{J}f$ $=\sup\{J(h)\,|\,h\in\mathscr{F},\,h\leq f\},$ (15)
$\bar{J}^{1}g$ $=\sup\{J^{-1}(g^{\prime})\,|\,g^{\prime}\in\mathcal{G},\,g^{\prime}\leq g\}.$ (16)
Clearly their restrictions to $\mathscr{F}$ and $\mathcal{G}$ are equal to $J$ and $J^{-1}$ respectively. Moreover $\bar{J}$ and $\bar{J}^{1}$ are order preserving. Fix $f\in\overline{\mathscr{F}}^{\sup}$, and set $\tilde{f}=\bar{J}^{1}\bar{J}f$. We want to show that $f=\tilde{f}$. We have that
@@
Let us show that $\tilde{f}\leq f$. Fix $h\in\mathscr{F}$ such that $J(h)\leq\sup_{h^{\prime}\leq f}J(h^{\prime})$. Set $g=J(h)$ and $g^{\prime}=\sup_{h^{\prime}\leq f}J(h^{\prime})$. Since $\mathcal{G}$ is finitely sup-stable, we can take $g^{\prime}=\lim\uparrow g_{\alpha}$ with each $g_{\alpha}$ a suprema of a finite number of functions $J(h^{\prime})$ satisfying $h^{\prime}\leq f$.color=blue!10,size=] PCA: détailler ? Hence by (12), for all $\epsilon>0$, there exists $g_{\alpha_{\epsilon}}$ such that $g\leq g_{\alpha_{\epsilon}}+\epsilon$, and $g_{\alpha_{\epsilon}}=\sup_{h_{i}^{\prime}\leq f,\,i\in\{1,\dots N\}}J(h_{i}^{\prime})$. Notice that in 7, (9) still holds in our setting for finite $\mathscr{A}$. Hence $J^{-1}$ commutes with a finite number of suprema. So we obtain $h\leq\sup_{i}J^{-1}(J(h^{\prime}_{i})+\epsilon)$. Applying (13), since $J(h^{\prime}_{i})\leq g^{\prime}$, for all $y\in\mathscr{Y}$, it holds $h(y)\leq\sup_{i}h^{\prime}_{i}(y)+v_{y}(\epsilon)\leq f(y)+v_{y}(\epsilon)$. Taking $\epsilon$ to 00, we get $h\leq f$.color=blue!10,size=] PCA: attention, si j’avais $\sup_{i}v_{y,i}(\epsilon)$, le nombre de $i$ dépendrait de $\epsilon$ ! C’est un peu pénible cette uniforme continuité.. Taking the supremum over $h$ shows that $\tilde{f}\leq f$.
To prove that $\tilde{f}\geq f$, we are going to show that, for any $f\in\overline{\mathscr{F}}^{\sup}$ and $g\in\overline{\mathcal{G}}^{\sup}$, we have that $g\geq\bar{J}f\implies\bar{J}^{1}g\geq f$. Take any $f,g$ satisfying $g\geq\bar{J}f$. Fix $f^{\prime}\in\mathscr{F}$ such that $f^{\prime}\leq f$, so $Jf^{\prime}\leq\bar{J}f\leq g$, whence $f^{\prime}\leq\bar{J}^{1}g$. Taking the supremum over $f^{\prime}$, we obtain $\bar{J}^{1}g\geq f$. Since $\bar{J}f\geq\bar{J}f$, the implication we have just shown entails that $\bar{J}^{1}\bar{J}f\geq f$. We thus obtained that $f=\bar{J}^{1}\bar{J}f$, so $\bar{J}$ is an order isomorphism.
For (14), just set $J(f):=J^{\prime}(f+f_{0})-g_{0}$ with inverse $J^{-1}(g)=(J^{\prime})^{-1}(g+g_{0})-f_{0}$. Take a sup-irreducible $f\in\mathscr{F}$, then, for all $g,h\in\mathscr{F}$, $f+f_{0}=\sup(g+f_{0},h+f_{0})\implies f=g\text{ or }f=h$, so $(f+f_{0})$ is sup-irreducible in $\mathscr{F}^{\prime}$. The other implications for irreducible functions are shown similarly. ∎
It was rightly mentioned in [AAM09, p670] that most results could be obtained either starting from order preserving or order reversing operators, without much difference. However we will focus on order preserving ones, as done in the convex case by [IRS15, p75]. Indeed order preserving operators have the advantage that the identity operator is always order preserving, whereas we will meet spaces with no anti-involution. Most of the following results stem from characterizing either the sup or inf-irreducible elements. The usefulness of focusing on such elements was already exemplified in [GK07, p5] for finite sets $\mathscr{X}$.